### v1

@@ 4.1 Superapproximation properties of element projector $\Pi$
$=\left|\frac{1}{2^{k+1+n}(k+1)!n!}\int_{\hat{K}}\partial_{s}^{k+1}\left[z\Big{(}x_{i+1/2}+\frac{1}{2}h_{i+1}s,y_{j-1/2}+\frac{1}{2}h_{j}t\Big{)}\right.\right.$
$\qquad\left.\left.-z\Big{(}x_{i-1/2}+\frac{1}{2}h_{i}s,y_{j-1/2}+\frac{1}{2}h_{j}t\Big{)}\right](s^{2}-1)^{k+1}\partial_{t}^{n}(t^{2}-1)^{n}\,\mathrm{d}s\,\mathrm{d}t\right|$
$=\frac{1}{2^{k+1+n}(k+1)!n!}\Big{(}\frac{1}{2}h_{i}\Big{)}^{k+1}\left|\int_{\hat{K}}\left[\int_{x_{i-1/2}}^{x_{i+1/2}}\partial_{x}^{k+2}z\left(x+\frac{1}{2}h_{i}s,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\mathrm{d}x\right]\right.$
$\hskip 28.45274pt\left.(s^{2}-1)^{k+1}\partial_{t}^{n}(t^{2}-1)^{n}\mathrm{d}s\,\mathrm{d}t\right|$
$\leq Ch_{i}^{k+1}\int_{-1}^{1}\int_{x_{i-1}}^{x_{i+1}}\left|\partial_{x}^{k+2}z\left(x,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\right|\,\mathrm{d}x\,\mathrm{d}t$
$=Ch_{i}^{k+1}h_{j}^{-1}\int_{y_{j-1}}^{y_{j}}\int_{x_{i-1}}^{x_{i+1}}|\partial_{x}^{k+2}z(x,y)|\,\mathrm{d}x\,\mathrm{d}y$
$\leq C\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}.$
Hence Cauchy-Schwarz and inverse inequalities yield
@@
$\leq Ch_{j}^{1/2}\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@
where $\left\|z\right\|_{L^{2}(\partial I_{i})}:=[z^{2}(x_{i}^{-})+z^{2}(x_{i-1}^{+})]^{1/2}$. Using this inequality, $\ell_{m}(x_{i})=\hat{\ell}_{m}(1)=1$, the definition of $\alpha_{x;m}^{ij}(x)$, a Cauchy-Schwarz inequality and $\left\|\ell_{n}\right\|_{J_{j}}=O(h_{j}^{1/2})$, we obtain
$\left|\sum_{m=k+2}^{\infty}\alpha^{ij}_{mn}\right|$ $=\left|\sum_{m=k+2}^{\infty}\alpha^{ij}_{mn}\ell_{m}(x_{i})\right|$
$=\left|(\alpha_{x;n}^{ij}-\pi^{k+1}\alpha_{x;n}^{ij})(x_{i})\right|$
$\leq\left\|\alpha_{x;n}^{ij}-\pi^{k+1}\alpha_{x;n}^{ij}\right\|_{L^{2}(\partial I_{i})}$
$\leq Ch_{i}^{k+3/2}\left\|\partial_{x}^{k+2}\alpha_{x;n}^{ij}\right\|_{I_{i}}$
@@
$=C\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}}.$ (4.10)
@@
$\left|\sum_{m=k+2}^{\infty}\alpha^{i+1,j}_{mn}(-1)^{m}\right|$ $\leq C\sqrt{\frac{h_{i+1}}{h_{j}}}\,h_{i+1}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{i+1,j}}$
$=C\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{i+1,j}}.$ (4.11)
Now the pairwise orthogonality of the rescaled Legendre polynomials and Cauchy-Schwarz and inverse inequalities yield
@@
$\leq Ch_{j}^{1/2}\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@ 4.2 Superapproximation properties of edge-element projectors
$\left|\sum_{n=k+1}^{\infty}\alpha^{ij}_{mn}\right|$ $\leq C\sqrt{\frac{h_{j}}{h_{i}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{ij}}.$ (4.16a)
@@
$\left|\sum_{n=k+1}^{\infty}\alpha^{i+1,j}_{mn}(-1)^{m}\right|$ $=\left|\sum_{n=k+1}^{\infty}\alpha^{i+1,j}_{mn}\right|\leq C\sqrt{\frac{h_{j}}{h_{i+1}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{i+1,j}}.$ (4.16b)
@@
Furthermore, imitating the derivation of (4.12) and recalling the inequalities (4.16), we obtain
@@
$\leq Ch_{j}^{1/2}\sqrt{\frac{h_{j}}{h_{i}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{ij}\cup K_{i+1,j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@ 5 Energy-norm and balanced-norm error estimates
Assume that $b$ is a positive constant and that Lemma 2.1 is valid for $m=k+1$. Let $\bm{w}=(u,p,q)$ be the true solution of the reaction-diffusion problem (1.1). Let $\bm{\mathbbm{W}}=(\mathbbm{U},\mathbbm{P},\mathbbm{Q})\in\mathcal{V}_{N}^{3}$ be the numerical solution of the LDG scheme (2.8) with layer-upwind flux (2.9) on the Shishkin mesh (2.6) with $\sigma\geq\hat{k}+1$, where $\hat{k}=k+1$ for even $k$ and $\hat{k}=k$ for odd $k$. Then there exists a constant $C>0$, which is independent of $\varepsilon$ and $N$, such that
@@
$T_{4}$ $=\left\langle\eta_{u},(\xi_{p})_{x}\right\rangle+\sum_{j=1}^{N}\sum_{i=1}^{N-1}\left\langle(\widehat{\eta_{u}})_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}},$
$T_{5}$ $=\left\langle\eta_{u},(\xi_{q})_{y}\right\rangle+\sum_{i=1}^{N}\sum_{j=1}^{N-1}\left\langle(\widehat{\eta_{u}})_{x,j},[\![\xi_{q}]\!]_{x,j}\right\rangle_{I_{i}},$
$T_{6}$ $=\left\langle\eta_{p},(\xi_{u})_{x}\right\rangle+\sum_{j=1}^{N}\left[\sum_{i=1}^{N-1}\left\langle(\widehat{\eta_{p}})_{i,y},[\![\xi_{u}]\!]_{i,y}\right\rangle_{J_{j}}+\left\langle(\eta_{p})_{0,y}^{+},(\xi_{u})_{0,y}^{+}\right\rangle_{J_{j}}-\left\langle(\eta_{p})_{N,y}^{-},(\xi_{u})_{N,y}^{-}\right\rangle_{J_{j}}\right],$
$T_{7}$ $=\left\langle\eta_{q},(\xi_{u})_{y}\right\rangle+\sum_{i=1}^{N}\left[\sum_{j=1}^{N-1}\left\langle(\widehat{\eta_{q}})_{x,j},[\![\xi_{u}]\!]_{x,j}\right\rangle_{I_{i}}+\left\langle(\eta_{q})_{x,0}^{+},(\xi_{u})_{x,0}^{+}\right\rangle_{I_{i}}-\left\langle(\eta_{q})_{x,N}^{-},(\xi_{u})_{x,N}^{-}\right\rangle_{I_{i}}\right].$
@@
$\left|\sum_{j=1}^{N}\left\langle(\eta_{p})_{N/4,y}^{+},(\xi_{u})^{-}_{N/4,y}\right\rangle_{J_{j}}\right|$
$\leq\left(\sum_{j=1}^{N}\left\|(\eta_{p})_{N/4,y}^{+}\right\|^{2}_{J_{j}}\right)^{1/2}\left(\sum_{j=1}^{N}\left\|(\xi_{u})_{N/4,y}^{-}\right\|^{2}_{J_{j}}\right)^{1/2}$
@@
$T_{4}=\sum_{j=1}^{N}\sum_{i=1}^{N}$ $\left[\left\langle u-\mathcal{I}u,(\xi_{p})_{x}\right\rangle_{K_{ij}}-\left\langle\widehat{(u-\mathcal{I}u)}_{i,y},(\xi_{p})^{-}_{i,y}\right\rangle_{J_{j}}\right.$
$\qquad\left.+\left\langle\widehat{(u-\mathcal{I}u)}_{i-1,y},(\xi_{p})^{+}_{i-1,y}\right\rangle_{J_{j}}\right]$
where $\widehat{(u-\mathcal{I}u)}_{i,y}(i=0,1,\dots,N)$ is specified by (2.9) with the special definitions $\widehat{(u-\mathcal{I}u)}_{0,y}=\widehat{(u-\mathcal{I}u)}_{N,y}=0$. Divide this sum into nine separate sums according to the different definitions of $\mathcal{I}$ in (3.3) over nine subregions, viz., $T_{4}:=\sum_{i=1}^{9}T_{4i}$, where
@@
$T_{44}$ $=\left\langle u-\Pi_{y}^{-}u,(\xi_{p})_{x}\right\rangle_{(\tau,1-\tau)\times(0,\tau)}+\sum_{j=1}^{N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{-}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}},$
$T_{45}$ $=\left\langle u-\Pi u,(\xi_{p})_{x}\right\rangle_{\Omega_{22}}+\sum_{j=N/4+1}^{3N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}},$
$T_{46}$ $=\left\langle u-\Pi_{y}^{+}u,(\xi_{p})_{x}\right\rangle_{(\tau,1-\tau)\times(1-\tau,1)}+\sum_{j=3N/4+1}^{N}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{+}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}},$
@@
A Cauchy-Schwarz inequality, inverse inequality, (3.6a) and (4.20a) show that
$|T_{44}|$ $\leq C\left[N\left\|u-\Pi_{y}^{-}u\right\|_{\Omega_{21}}\left\|\xi_{p}\right\|_{\Omega_{21}}+\left|\sum_{j=1}^{N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{-}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}\right|\right]$
$\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}}.$
@@
By the definition (3.1a) of the $L^{2}$-projector $\Pi$ and (4.13a), we get
@@
Then the bounds on the terms in (5.1) remain unchanged except for $T_{4}$ and $T_{5}$. In the subdivision of $T_{4}$, we will have a different expression for the term $T_{45}$:
$T_{45}=\left\langle u-\Pi_{b}u,(\xi_{p})_{x}\right\rangle_{\Omega_{22}}+\sum_{j=N/4+1}^{3N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{b}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}:=T_{45}^{1}+T_{45}^{2}.$
@@
Adding (5.15) and (5.18) gives the bound $T_{45}\leq CN^{-\hat{k}}\left\|\xi_{p}\right\|_{\Omega_{22}}$ and thus the same bound for $T_{4}$. We can bound $T_{5}$ in a similar way. Thus, we again obtain the error bound of Theorem 5.1.

### v3

@@ 4.1 Superapproximation properties of element projector $\Pi$
$=\frac{2k+3}{2^{k+2}(k+1)!}\cdot\frac{1}{\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\left|\int_{\hat{K}}\partial_{s}^{k+1}\left[z\Big{(}x_{i+1/2}+\frac{1}{2}h_{i+1}s,y_{j-1/2}+\frac{1}{2}h_{j}t\Big{)}\right.\right.$
$\qquad\left.\left.-z\Big{(}x_{i-1/2}+\frac{1}{2}h_{i}s,y_{j-1/2}+\frac{1}{2}h_{j}t\Big{)}\right](s^{2}-1)^{k+1}\hat{\ell}_{n}(t)\,\mathrm{d}s\,\mathrm{d}t\right|$
$=\frac{2k+3}{2^{k+2}(k+1)!}\cdot\frac{1}{\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\Big{(}\frac{1}{2}h_{i}\Big{)}^{k+1}\left|\int_{\hat{K}}\left[\int_{x_{i-1/2}}^{x_{i+1/2}}\partial_{x}^{k+2}z\left(x+\frac{1}{2}h_{i}s,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\mathrm{d}x\right]\right.$
$\hskip 28.45274pt\left.(s^{2}-1)^{k+1}\hat{\ell}_{n}(t)\mathrm{d}s\,\mathrm{d}t\right|$
$\leq Ch_{i}^{k+1}\frac{1}{\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\int_{-1}^{1}\int_{x_{i-1}}^{x_{i+1}}\left|\partial_{x}^{k+2}z\left(x,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\right||\hat{\ell}_{n}(t)|\,\mathrm{d}x\,\mathrm{d}t$
$\leq Ch_{i}^{k+1}\frac{1}{\left\|\hat{\ell}_{n}\right\|^{2}_{\hat{I}}}\left[\int_{-1}^{1}\int_{x_{i-1}}^{x_{i+1}}\left|\partial_{x}^{k+2}z\left(x,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\right|^{2}\,\mathrm{d}x\,\mathrm{d}t\right]^{1/2}\cdot\left(h_{i}^{1/2}\left\|\hat{\ell}_{n}\right\|_{\hat{I}}\right)$
$=C\sqrt{2n+1}\cdot h_{i}^{k+3/2}h_{j}^{-1/2}\left[\int_{y_{j-1}}^{y_{j}}\int_{x_{i-1}}^{x_{i+1}}\left|\partial_{x}^{k+2}z\left(x,y_{j-1/2}+\frac{1}{2}h_{j}t\right)\right|^{2}\,\mathrm{d}x\,\mathrm{d}y\right]^{1/2}$
$=C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}.$
Hence Cauchy-Schwarz and inverse inequalities and (4.2) yield
@@
$\leq C\max_{0\leq n\leq k}\Big{(}|\alpha^{ij}_{k+1,n}-\alpha^{i+1,j}_{k+1,n}|\Big{)}\left\|\ell_{n}\right\|_{J_{j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
$\leq C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}\sqrt{\frac{h_{j}}{2n+1}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@
where $\left\|z\right\|_{L^{2}(\partial I_{i})}:=[z^{2}(x_{i}^{-})+z^{2}(x_{i-1}^{+})]^{1/2}$. Using this inequality, $\ell_{m}(x_{i})=\hat{\ell}_{m}(1)=1$, the definition of $\alpha_{x;m}^{ij}(x)$, a Cauchy-Schwarz inequality and (4.2), we obtain
$\left|\sum_{m=k+2}^{\infty}\alpha^{ij}_{mn}\right|$ $=\left|\sum_{m=k+2}^{\infty}\alpha^{ij}_{mn}\ell_{m}(x_{i})\right|=\left|(\alpha_{x;n}^{ij}-\pi^{k+1}\alpha_{x;n}^{ij})(x_{i})\right|$
$\leq\left\|\alpha_{x;n}^{ij}-\pi^{k+1}\alpha_{x;n}^{ij}\right\|_{L^{2}(\partial I_{i})}\leq Ch_{i}^{k+3/2}\left\|\partial_{x}^{k+2}\alpha_{x;n}^{ij}\right\|_{I_{i}}$
@@
$=C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}}.$ (4.11)
@@
$\left|\sum_{m=k+2}^{\infty}\alpha^{i+1,j}_{mn}(-1)^{m}\right|$ $\leq C\sqrt{2n+1}\sqrt{\frac{h_{i+1}}{h_{j}}}\,h_{i+1}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{i+1,j}}$
$=C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}\,h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{i+1,j}}.$ (4.12)
Now the pairwise orthogonality of the rescaled Legendre polynomials, (4.2) and Cauchy-Schwarz and inverse inequalities yield
@@
$\leq C\sqrt{2n+1}\sqrt{\frac{h_{i}}{h_{j}}}h_{i}^{k+1}\left\|\partial_{x}^{k+2}z\right\|_{K_{ij}\cup K_{i+1,j}}\sqrt{\frac{h_{j}}{2n+1}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@ 4.2 Superapproximation properties of edge-element projectors
$\left|\sum_{n=k+1}^{\infty}\alpha^{ij}_{mn}\right|$ $\leq C\sqrt{2m+1}\sqrt{\frac{h_{j}}{h_{i}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{ij}}.$ (4.17a)
@@
$\left|\sum_{n=k+1}^{\infty}\alpha^{i+1,j}_{mn}(-1)^{m}\right|$ $=\left|\sum_{n=k+1}^{\infty}\alpha^{i+1,j}_{mn}\right|\leq C\sqrt{2m+1}\sqrt{\frac{h_{j}}{h_{i+1}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{i+1,j}}.$ (4.17b)
@@
Furthermore, imitating the derivation of (4.13), (4.2) and recalling the inequalities (4.17), we obtain
@@
$\leq C\max_{0\leq m\leq k}\left[\left|\sum_{n=k+1}^{\infty}\alpha^{ij}_{mn}\right|+\left|\sum_{n=k+1}^{\infty}\alpha^{i+1,j}_{mn}(-1)^{m}\right|\right]\left\|\ell_{k}\right\|_{J_{j}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
$\leq C\sqrt{2k+1}\sqrt{\frac{h_{j}}{h_{i}}}h_{j}^{k}\left\|\partial_{y}^{k+1}z\right\|_{K_{ij}\cup K_{i+1,j}}\sqrt{\frac{h_{j}}{2k+1}}h_{i}^{-1/2}\left\|\mathbbm{v}\right\|_{K_{ij}\cup K_{i+1,j}}$
@@ 5 Energy-norm and balanced-norm error estimates
Assume that $b$ is a positive constant and that Lemma 2.1 is valid for $m=k+1$. Let $\bm{w}=(u,p,q)=(u,\varepsilon u_{x},\varepsilon u_{y})$ be the true solution of the reaction-diffusion problem (1.1). Let $\bm{\mathbbm{W}}=(\mathbbm{U},\mathbbm{P},\mathbbm{Q})\in\mathcal{V}_{N}^{3}$ be the numerical solution of the LDG scheme (2.8) with layer-upwind flux (2.9) on the Shishkin mesh (2.6) with $\sigma\geq\hat{k}+1$, where $\hat{k}=k+1$ for even $k$ and $\hat{k}=k$ for odd $k$. Then there exists a constant $C>0$, which is independent of $\varepsilon$ and $N$, such that
@@
$T_{4}$ $=\left\langle\eta_{u},(\xi_{p})_{x}\right\rangle+\sum_{j=1}^{N}\sum_{i=1}^{N-1}\left\langle(\widehat{\eta_{u}})_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}},\;T_{5}=\left\langle\eta_{u},(\xi_{q})_{y}\right\rangle+\sum_{i=1}^{N}\sum_{j=1}^{N-1}\left\langle(\widehat{\eta_{u}})_{x,j},[\![\xi_{q}]\!]_{x,j}\right\rangle_{I_{i}},$
$T_{6}$ $=\left\langle\eta_{p},(\xi_{u})_{x}\right\rangle+\sum_{j=1}^{N}\sum_{i=0}^{N}\left\langle(\widehat{\eta_{p}})_{i,y},[\![\xi_{u}]\!]_{i,y}\right\rangle_{J_{j}},\;T_{7}=\left\langle\eta_{q},(\xi_{u})_{y}\right\rangle+\sum_{i=1}^{N}\sum_{j=0}^{N}\left\langle(\widehat{\eta_{q}})_{x,j},[\![\xi_{u}]\!]_{x,j}\right\rangle_{I_{i}}.$
@@
$\left|\sum_{j=1}^{N}\left\langle(\eta_{p})_{N/4,y}^{+},(\xi_{u})^{-}_{N/4,y}\right\rangle_{J_{j}}\right|$ $\leq\left(\sum_{j=1}^{N}\left\|(\eta_{p})_{N/4,y}^{+}\right\|^{2}_{J_{j}}\right)^{1/2}\left(\sum_{j=1}^{N}\left\|(\xi_{u})_{N/4,y}^{-}\right\|^{2}_{J_{j}}\right)^{1/2}$
@@
$T_{4}=\sum_{j=1}^{N}\sum_{i=1}^{N}\Big{[}\left\langle\eta_{u},(\xi_{p})_{x}\right\rangle_{K_{ij}}-\left\langle(\widehat{\eta_{u}})_{i,y},(\xi_{p})^{-}_{i,y}\right\rangle_{J_{j}}+\left\langle(\widehat{\eta_{u}})_{i-1,y},(\xi_{p})^{+}_{i-1,y}\right\rangle_{J_{j}}\Big{]}$
where $(\widehat{\eta_{u}})_{i,y}(i=0,1,\dots,N)$ is specified by (2.9) with the special definitions $(\widehat{\eta_{u}})_{0,y}=(\widehat{\eta_{u}})_{N,y}=0$. Divide this sum into nine separate sums according to the different definitions of $\mathcal{I}$ in (3.3) over nine subregions, viz., $T_{4}:=\sum_{i=1}^{9}T_{4i}$, where
@@
$T_{44}$ $=\left\langle u-\Pi_{y}^{-}u,(\xi_{p})_{x}\right\rangle_{\Omega_{21}}+\sum_{j=1}^{N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{-}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}$
$\quad+\sum_{j=1}^{N/4}\left\langle(u-\Pi^{-}_{xy}u)^{-}_{N/4,y},(\xi_{p})^{+}_{N/4,y}\right\rangle_{J_{j}}-\sum_{j=1}^{N/4}\left\langle(u-\Pi^{\pm}_{xy}u)^{+}_{3N/4,y},(\xi_{p})^{-}_{3N/4+1,y}\right\rangle_{J_{j}},$
$T_{45}$ $=\left\langle u-\Pi u,(\xi_{p})_{x}\right\rangle_{\Omega_{22}}+\sum_{j=N/4+1}^{3N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}$
$\quad+\sum_{j=N/4+1}^{3N/4}\left\langle(u-\Pi^{-}_{x}u)^{-}_{N/4,y},(\xi_{p})^{+}_{N/4,y}\right\rangle_{J_{j}}-\sum_{j=N/4+1}^{3N/4}\left\langle(u-\Pi^{+}_{x}u)^{+}_{3N/4,y},(\xi_{p})^{-}_{3N/4+1,y}\right\rangle_{J_{j}},$
$T_{46}$ $=\left\langle u-\Pi_{y}^{+}u,(\xi_{p})_{x}\right\rangle_{\Omega_{23}}+\sum_{j=3N/4+1}^{N}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{+}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}$
$\quad+\sum_{j=3N/4+1}^{N}\left\langle(u-\Pi^{\mp}_{xy}u)^{-}_{N/4,y},(\xi_{p})^{+}_{N/4,y}\right\rangle_{J_{j}}-\sum_{j=3N/4+1}^{N}\left\langle(u-\Pi^{+}_{xy}u)^{+}_{3N/4,y},(\xi_{p})^{-}_{3N/4+1,y}\right\rangle_{J_{j}},$
@@
A Cauchy-Schwarz inequality, an inverse inequality and (3.6a) show that
$|\left\langle u-\Pi_{y}^{-}u,(\xi_{p})_{x}\right\rangle_{\Omega_{21}}|$ $\leq CN\left\|u-\Pi_{y}^{-}u\right\|_{\Omega_{21}}\left\|\xi_{p}\right\|_{\Omega_{21}}$
$\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}}.$ (5.14)
The inequality (4.21a) gives
$\left|\sum_{j=1}^{N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{y}^{-}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}\right|\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}}.$ (5.15)
Using the Cauchy-Schwarz and inverse inequalities and (3.6b) yields
$\left|\sum_{j=1}^{N/4}\left\langle(u-\Pi^{-}_{xy}u)^{-}_{N/4,y},(\xi_{p})^{+}_{N/4,y}\right\rangle_{J_{j}}\right|$ $\leq\sum_{j=1}^{N/4}\left\|(u-\Pi^{-}_{xy}u)^{-}_{N/4,y}\right\|_{J_{j}}\left\|(\xi_{p})^{+}_{N/4,y}\right\|_{J_{j}}$
$\leq C\sum_{j=1}^{N/4}h_{j}^{1/2}\left\|\eta_{u}\right\|_{L^{\infty}(K_{N/4,j})}N^{1/2}\left\|\xi_{p}\right\|_{K_{N/4+1,j}}$
$\leq C\tau^{1/2}N^{1/2}\left\|\eta_{u}\right\|_{L^{\infty}(\Omega_{11})}\left\|\xi_{p}\right\|_{\Omega_{21}}$
$\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}},$ (5.16)
where $N^{-1/2}(\ln N)^{1/2}\leq C$ was used for $N\geq 4$. Similarly,
$\left|\sum_{j=1}^{N/4}\left\langle(u-\Pi^{\pm}_{xy}u)^{+}_{3N/4,y},(\xi_{p})^{-}_{3N/4+1,y}\right\rangle_{J_{j}}\right|\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}}.$ (5.17)
Combining (5.14)–(5.17) yields
$|T_{44}|\leq C\varepsilon^{1/4}N^{-k}(\ln N)^{k+1}\left\|\xi_{p}\right\|_{\Omega_{21}}.$
@@
Appealing first to the definition (3.1) of the $L^{2}$-projector $\Pi$ and the Gauss-Radau projectors $\Pi^{-}_{x}$ and $\Pi^{+}_{x}$, we use (4.14a) to get
@@
Our LDG method does not employ any penalty terms. If one uses a numerical flux with penalty in the LDG equations (2.8), e.g., ^$\widehat{\mathbbm{P}}^{\star}_{i,y}=\widehat{\mathbbm{P}}_{i,y}+\lambda_{i}[\![\mathbbm{U}]\!]_{i,y}$ and ^$\widehat{\mathbbm{Q}}^{\star}_{x,j}=\widehat{\mathbbm{Q}}_{x,j}+\mu_{j}[\![\mathbbm{U}]\!]_{x,j}$ for $i,j=0,1,\dots,N$, where ^$\widehat{\mathbbm{P}}_{i,y}$ and ^$\widehat{\mathbbm{Q}}_{x,j}$ are defined in (2.9) and the penalty parameters $\lambda_{i}=\mu_{j}=\varepsilon^{1/2}$ for $i,j=0,1,\dots,N$, then modify our error analysis accordingly, one obtains an optimal-order error estimate in the corresponding energy and balanced norms. But the convergence rate is heavily influenced by the error from the boundary penalty terms; consequently, the theoretical result for the $H^{1}$-error is only a suboptimal convergence rate $O((N^{-1}\ln N)^{k+1/2})$, which is a half-order inferior to the numerical results.
Remark 5.7.
In our analysis we considered the standard Shishkin mesh, whose coarse and fine components are each piecewise uniform. If instead a randomly perturbed quasi-uniform mesh is used in $\Omega_{22}$, then a slightly worse bound can be derived in Lemmas 4.1 and 4.3, and finally the term $\varepsilon^{1/4}N^{-\hat{k}}$ that appears in the balanced-norm error bound will be changed to $\varepsilon^{1/4}N^{-k}$. Because of the positive power of the small parameter, the final convergence rate is only slightly affected. Furthermore, our analysis is applicable to other layer-adapted meshes such as Bakhvalov-type meshes which are not locally uniform inside the layer regions, but the convergence result will be significantly different. The extension of our methodology to other types of layer-adapted meshes deserves further investigation and we shall explore this in future work, as we state in Section 7.
Remark 5.8.
@@
Then the bounds on the terms in (5.1) remain unchanged except for $T_{4}$ and $T_{5}$. In the subdivision of $T_{4}$, we only have a different expression for the former two terms included in $T_{45}$, denoted by
$T_{45}^{1}:=\left\langle u-\Pi_{b}u,(\xi_{p})_{x}\right\rangle_{\Omega_{22}}\text{ and }T_{45}^{2}:=\sum_{j=N/4+1}^{3N/4}\sum_{i=N/4+1}^{3N/4-1}\left\langle\{\!\{u-\Pi_{b}u\}\!\}_{i,y},[\![\xi_{p}]\!]_{i,y}\right\rangle_{J_{j}}.$
@@
Adding (5.19), (5.22) and the existing bounds for the remaining terms of $T_{45}$ gives $T_{45}\leq CN^{-\hat{k}}\left\|\xi_{p}\right\|_{\Omega_{22}}$ and thus the same bound for $T_{4}$. We can bound $T_{5}$ in a similar way. Thus, we again obtain the error bound of Theorem 5.1.